OpenStudy (anonymous):

Which inequality represents the solutions to 2a - 20<28? A. a < 4 B. a > 4 C. a < 24 D. a > 24

OpenStudy (dmezzullo):

Ok so first off u wanna get the like terms together.

OpenStudy (dmezzullo):

So You would have to add 20 on both sides. then wat would u have?

OpenStudy (dmezzullo):

@dinosour1480

OpenStudy (anonymous):

well 20 + 20=40

OpenStudy (dmezzullo):

No, U wanna add 20 to both sides so it would be 2a<48

OpenStudy (anonymous):

oh sorry

OpenStudy (dmezzullo):

So then what would u get wen u divided 2a<48?

OpenStudy (anonymous):

divide by 20?

OpenStudy (dmezzullo):

Divide by 2a

OpenStudy (anonymous):

2*2=1

OpenStudy (anonymous):

48*2=24

OpenStudy (anonymous):

so it's c

OpenStudy (dmezzullo):

Ok so put that together and choice

OpenStudy (dmezzullo):

choose*

OpenStudy (dmezzullo):

Why do you think its C?

OpenStudy (dmezzullo):

@dinosour1480

OpenStudy (anonymous):

because 24 is greater than 1

OpenStudy (dmezzullo):

Wait ya ur right

OpenStudy (dmezzullo):

sorry i confused myself

OpenStudy (anonymous):

thanks for the help mezz